OpenStudy (lawliet):

One of the solutions to x2 − 2x − 15 = 0 is x = −3. What is the other solution? x = −5 x = −1 x = 1 x = 5

OpenStudy (rawritskai):

hmm how would i explain this properly

OpenStudy (lawliet):

im not good at math and thats all i have left to do

OpenStudy (rawritskai):

i know the answer but umm idk how to explain it in a good way :/

OpenStudy (lawliet):

i chose the answer B because .... life \

OpenStudy (lawliet):

some parts of algebvra im good at others like this im not

OpenStudy (lawliet):

nor is my spelling good

OpenStudy (rawritskai):

okay thats not correct try again :) which seems more reasonable

OpenStudy (rawritskai):

and its fine lol

OpenStudy (lawliet):

i dont know i didnt have an algebra class for about three and a half months freshman year

OpenStudy (lawliet):

because i wasnt in school

OpenStudy (lawliet):

lemme try it again

OpenStudy (lawliet):

one of the answers is 5

OpenStudy (lawliet):

or d

OpenStudy (rawritskai):

hmmmm okay this is how i can easily explain it x = [−b ± √(b2 − 4ac)]/2a a = 1, b = 2, and c = −15. Thus: x = [−2 ± √(22 − 4*1*{−15})]/2 x = [−2 ± √(4 + 60)]/2 x = [−2 ± √(64)]/2 x = [−2 ± 8]/2 The two solutions are: x = −10/2 and x = +6/2 x = −5 and x = 3

OpenStudy (rawritskai):

yes correct :)

OpenStudy (lawliet):

you get a medal miss lady

OpenStudy (rawritskai):

thanks lol, you dont remeber me do u

OpenStudy (lawliet):

i do

OpenStudy (lawliet):

help with this one too please

OpenStudy (lawliet):

Which represents the solution(s) of the system of equations, y = x2 – x + 1 and y = x? Determine the solution set by graphing. (1, 1) (0, 1) and (1, 1) (0, 0) and (1, 1) no solutions

OpenStudy (lawliet):

just to make sure im doing this right i chose a or (1,1)

OpenStudy (rawritskai):

mmkkk lemme see give me a sec to look it over

OpenStudy (lawliet):

mkay

OpenStudy (rawritskai):

I'd have to say its B since it also levels out at 0,1 on the graph as well

OpenStudy (lawliet):

ill look at it again thank you pm me if you want

OpenStudy (rawritskai):

mmhhm np and okay